Klapperich Lab:Notebook/Lab Meeting Notes/2010/04/15

From OpenWetWare
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

15 April 2010 Lab Meeting

Jake Presentation.

‡ Announcements

‡ Flu R01:Integration

  • meetings every other M 1-3pm. 705.

  • for PCR just do more reactions, pool and etoh precipitate.
    • For HDA, we need to explore cloning.

† HDA (Lead: Jaephil, Team:Sonali)

  • IBC Approval or 720 to work with real C.diff samples.
  • Patient Samples - dilution experiments (MM/SH)
  • Start planning R01 for Submission on 10/5/10. MM, CMK.
  • HDA mix in dried form - Expecting it from BH -- ON HOLD.
  • warm start here (Give some to Lisa)
  • QQ preliminary chip work
  • Sequence all output
  • 2nd HDA design - "microSPE" + "Flu primer" + "HDA in a cup" + "Detection kit from Biohelix"

† Sample Concentration (Lead: Jane, Team: Jaephil)

  • Straight channel adhesion test PCR result expected. From 4/6 result, PDMS 4mm wide channel doesn't show decrease in output RNA over time
  • Straight 1mm and 4mm PDMS-Teflon channel adhesion test pending PCR result.
  • Try collection vs dilutions, collection vs time, need binding constant to quantify results.
  • Samples expected from Fauchet group. Contact today. ??? - CMK will check again.

† SPE Column Optimization for DNA/RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)

  • (Almost done) Jessie will run chips with serial virus dilutions from 1:1 to 1:1,000,000 to re-do previous dilution experimnent with WHO SYBR green assay on chip.

† Integrated chip for flu(SPE+RT+PCR) (Lead: Qingqing)

  • Real copy number of previous human samples. So QQ is working with the higher concentration samples. Range is 10^6-10^8 of copy number/ml.
  • REVISIT the % recovery number.
  • QQ tried: more enzyme, Tween 20, BSA amounts.
  • integrated flu chip paper has been send to cathie, wait for revision.

  • PCR1 will combine with PCR2, waitting for revision (CMK)

‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )

  • See email notes.

‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)

  • New meeting time, T, 9:30 am every other week. MEETINGS CANCELLED NEXT WEEK (4/20).

‡ Agilent Automated Sample Preparation (Lead: Alex)

  • Paper on HotDog - with CMK/ASB

‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)

  • Evap paper - Additional experiment for Fig 3. JD says will send by Friday (4/16).

‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
* Cathie will submit paper.

‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)

  • Sample collection ongoing.

‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME))

  • .
  • Frank: Two- and three-week storage experiments completed. Four-week to be done next Wed. Longer-term (2, 3, 4-month) in storage. +/- carrier RNA not yet completed. Currently writing up final report.
  • Sean: First set of data is in (for the 700 nm silica straws). Today (4/15) I will finish the straws for the 150 nm silica and the SPE batters for the 14 nm silica. Tomorrow I plan on running and PCR'ing samples for the 150 nm silica straws as well as making the straws for the 14 nm silica. On Saturday I will run and PCR samples for the last set of straws.
  • Mark: as of 4/15/10 up to 30% recovery from 10 ng input sample with water elution. Also up to 13% recovery with 100ng/straw. 1-week trial completed as of 3/24. Results peak at 20% (-70C) and decrease with temperature. 2-month and 1-month trials complete, resulting in further degradation than in the 1-week trial. 4- and 6-month trials still in storage.

‡ IIH Senior Project.

  • ELISA done, Talking to Phil Allen about proper imaging
  • Writing report